how much water should be evaporated from 24 oz of a 4% solution to make a 6% salt solution
Answer by RAY100(1637)

Let x = amt of water evaporated,,,water is 0 % salt
.
(.04) 24 - (0) x = (24-x) (.06)
.
.96 = 1.44 - .06x
.
.06 x = 1.44 - .96 = .48
.
x=8 oz
.
.
check,,.04*24 = 16*.06 = .96,,,,,,,ok
.